# On the randomized Euler scheme for SDEs with integral-form drift

### v1

@@ 4. Error of the randomized Euler scheme
In this section, we introduce a randomized Euler scheme for the strong approximation of solutions of (1) and investigate its $p$-th moments, informational cost, and error upper bounds.
If one can access information about $a$ directly, which is possible if the explicit formula of $a$ is assumed to be known, the standard Euler-Maruyama scheme can be leveraged to approximate the value of $X(T)$. The scheme is defined as
@@
Its $L^{p}(\Omega)$-error, for $p\geq 2,$ is proportional to $\mathcal{O}(n^{-1/2})$; see Theorem 10.2.2 in [14]. On the other hand, the informational cost of the standard algorithm is always $\mathcal{O}(n).$
Nevertheless, the information about $a$ may be unknown whereas one knows the exact formula of $H.$ In the next part, we introduce a randomized Euler algorithm that can access the information about function $H$ instead of $a.$ Let $n\in\mathbb{N},(M_{k})_{k=1}^{n}\subset\mathbb{N},h=T/n$ and $t_{k}=kh$ for $k=0,1.\ldots,n.$ randomized Euler scheme is defined as
$\begin{cases}X_{n,M}^{RE}(0)=\eta,\\ X_{n,M}^{RE}(t_{k+1})=X_{n,M}^{RE}(t_{k})+\frac{h}{M_{k+1}}\sum\limits_{j=1}^{M_{k+1}}H(\xi_{j}^{k+1},X_{n,M}^{RE}(t_{k}))+b(X_{n,M}^{RE}(t_{k}))\Delta W_{k},\\ \quad k=0,1,\ldots,n-1\end{cases}$ (7)
where $(\xi_{j}^{k})_{j,k}$ is $\xi$-distributed i.i.d sequence and $\sigma\Big{(}\bigcup\limits_{k\geq 0,j\geq 1}\sigma(\xi_{j}^{k})\Big{)}\perp\!\!\!\perp\sigma\Big{(}\bigcup\limits_{t\geq 0}\mathcal{F}_{t}^{W}\Big{)}$ which means the randomization is independent of driving Wiener process. The continuous version of the randomized Euler scheme is defined as
$\tilde{X}_{n,M}^{RE}(t)=\eta+\int\limits_{0}^{t}\sum\limits_{k=0}^{n-1}\frac{1}{M_{k+1}}\sum\limits_{j=1}^{M_{k+1}}H(\xi_{j}^{k+1},\tilde{X}_{n,M}^{RE}(t_{k}))\mathds{1}_{(t_{k},t_{k+1}]}(s)\,{\mathrm{d}}s+\int\limits_{0}^{t}\sum\limits_{k=0}^{n-1}b(\tilde{X}_{n,M}^{RE}(t_{k}))\mathds{1}_{(t_{k},t_{k+1}]}(s)\,{\mathrm{d}}W(s)$
for $t\in[0,T].$ The continuous version of standard Euler algorithm $\tilde{X}_{n}^{E}$ is defined analogously. Depending on the context, $M$ in the subscript may either denote a sequence $(M_{k})_{k=1}^{n}$ or $M:=M_{1}\in\mathbb{N}$ if the sequence is constant. From mathematical induction, one obtains that
@@
for all $n$ and $M.$ Furthermore, from fact 2, the randomized Euler scheme takes finite values.
@@
There exists constant $C\in[0,+\infty)$ depending only on $\eta,T$ and parameters of the class $\mathcal{F}(L,K,p),$ such that for all $(a,b)\in\mathcal{F}(L,K,P),M,n\in\mathbb{N}$ the following condition holds
@@
Next, we define the complexity of the algorithm in terms of the information-based complexity framework; see [25]. The algorithm evaluates the value of $W$ at $n$ points, $b(\cdot)$ at $n$ points and $H(\cdot,\cdot)$ at $\sum\limits_{k=0}^{n-1}M_{k+1}$ points. Thus, $\operatorname{cost}(X_{n,M}^{RE})=2n+\sum\limits_{k=0}^{n-1}M_{k+1},$ and if $M_{k}=M$ for all $k=1,\ldots,n,$ then $\operatorname{cost}(X_{n,M}^{RE})=2n+nM=\mathcal{O}(nM).$ One may notice that the informational cost differs from the informational cost of the standard (randomized) Euler scheme.
@@
There exists constant $C\in[0,+\infty)$ depending only on $\eta,T$ and parameters of the class $\mathcal{F}(L,K,p),$ such that for all $(a,b)\in\mathcal{F}(L,K,p)$ and $n,M\in\mathbb{N}$ the following inequality holds
@@
In this proof, we leverage a continuous version of Euler algorithms. First, note that
@@
$\begin{split}A_{1,n}^{M}(t)&\leq\|L\|_{L^{1}}^{p}\sum\limits_{k=0}^{n-1}\int\limits_{0}^{t}\mathbb{E}\Big{\|}\tilde{X}_{n,M}^{RE}(t_{k})-\tilde{X}_{n}^{E}(t_{k})\Big{\|}^{p}\mathds{1}_{(t_{k},t_{k+1}]}(s)\,{\mathrm{d}}s\\ &\leq\|L\|_{L^{1}}^{p}\int\limits_{0}^{t}\sup\limits_{0\leq u\leq s}\mathbb{E}\Big{\|}\tilde{X}_{n,M}^{RE}(u)-\tilde{X}_{n}^{E}(u)\Big{\|}^{p}\,{\mathrm{d}}s\end{split}$
@@
$\begin{split}\mathbb{E}\|B_{n}^{M}(t)\|^{p}&\leq C_{p}\mathbb{E}\Big{(}\int\limits_{0}^{t}\sum_{k=0}^{n-1}\|b(\tilde{X}_{n,M}^{RE}(t_{k}))-b(\tilde{X}_{n}^{E}(t_{k}))\|^{2}\mathds{1}_{(t_{k},t_{k+1}]}(s)\,{\mathrm{d}}s\Big{)}^{p/2}\\ &\leq C_{p}T^{p/2-1}\int\limits_{0}^{t}\sum_{k=0}^{n-1}\mathbb{E}\Big{\|}b(\tilde{X}_{n,M}^{RE}(t_{k}))-b(\tilde{X}_{n}^{E}(t_{k}))\Big{\|}^{p}\mathds{1}_{(t_{k},t_{k+1}]}(s)\,{\mathrm{d}}s\\ &\leq C_{p}T^{p/2-1}K^{p}\int\limits_{0}^{t}\sup\limits_{0\leq u\leq s}\mathbb{E}\Big{\|}\tilde{X}_{n,M}^{RE}(u)-\tilde{X}_{n}^{E}(u)\Big{\|}^{p}\,{\mathrm{d}}s.\\ \end{split}$
@@
Since function function $[0,T]\ni t\mapsto\sup\limits_{0\leq u\leq t}\mathbb{E}\|\tilde{X}_{n,M}^{RE}(u)-\tilde{X}_{n}^{E}(u)\|^{p}$ is Borel-measurable (as a non-decreasing function) and bounded (by fact 2), applying the Gronwall’s lemma yields
@@
From Minkowski inequality, Theorem 10.2.2 in [14] and lemma (2), one obtains that
@@ 5. Numerical experiments
In this section, we compare the obtained theoretical results with the outputs of performed simulations. In subsection 5.1, we estimate $L^{2}(\Omega)$ error of the randomized Euler algorithm, and compare it with its informational cost. In subsection 5.2, we inspect randomized Euler algorithm trajectories in the optimization problem of finding local minima. Finally, in subsection 5.3 we provide implementation details in Python and CUDA C.
@@ 5.1. Upper error bound vs informational cost
$b(t,x_{1},x_{2})=\begin{pmatrix}0.16x_{1}&0.24x_{2}\\ 0.24x_{1}&0.32x_{2}\\ \end{pmatrix}.$
@@
The figure 1 shows results obtained via numerical experiments for $\varepsilon_{6400}(X^{RE}_{n,n}(1))$ and various values of $n.$ The implementation utilizes CUDA C architecture using $100$ blocks and $32$ threads per block. Consequently, we obtain a strong linear correlation between estimated error and cost. The slope coefficient is equal to $-3.82$ which is seemingly close to the expected slope of order $-4.$
@@ 5.3. Implementation details
In this subsection, we provide implementation details in Python and CUDA C.
1def example():
2 # (1)
3 return sdes.SDE(
4 # (2)
5 drift=sdes.Function(
6 args=("int d", "t", "type* x"),
7 body=f"0.4*(d+1)*pow(t,2)*sin(x[1-d])"
8 ),
9 # (3)
10 diffusion=sdes.Function(
11 args=("int d", "int k", "t", "type* x"),
12 body="(d+k+1)*0.08*x[k]"
13 ),
14 DRIFT_DISTR="normal()", # (4)
15 precision=np.float64 # (5)
16 )
Listing 1: SDE example
@@ 14 DRIFT_DISTR="normal()", # (4)
    Create and return example instance of the SDE class.
@@
    Pass drift definition to the constructor.
@@
    Pass diffusion definition to the constructor.
  4. (4)
    Specify the distribution of the randomized drift.
  5. (5)
    Specify floating point precision.
1__device__ void sample_from_coupled_X(int ML, int Ml, int nL, int nl, int dW_dim, FP* x0, FP T, FP* XL, FP* Xl, int X_dim, curandState_t* state) {
@@ 21 FP t, dW;
38 ksi[j] = drift_inverse_cdf<FP>((FP) DRIFT_DISTR); // (7)
@@
1template <typename type>
2__device__ void euler_single_step(type t, type* X_prev, int X_dim, type h, type* ksi, int ksi_num, type* dW, int dW_dim, type* X) {
3 // (1)
4 for (int d = 0; d < X_dim; d++){
5 // (2)
6 for (int j = 0; j < ksi_num; j++) {
7 X[d] += func_a<type>(d, ksi[j], X_prev) * (h/ksi_num);
8 }
9 // (3)
10 for (int k = 0; k < dW_dim; k++) {
11 X[d] += func_b<type>(d, k + 1, t, X_prev) * dW[k];
12 }
13 }
14}
Listing 3: Numerical scheme implementation
  1. (1)
    Traverse through the equations.
  2. (2)
    Add drift component.
  3. (3)
    Add diffusion component.
@@ 6. Conclusions
We have investigated the error of the randomized Euler scheme in the case when the drift of underlying SDE has a special integral form. We conducted numerical experiments using GPU architecture. The numerical experiments show that the randomized Euler scheme is a suitable method for approximating trajectories of SDEs 1 and, on the other hand, it can be used as an optimization method (due to its connection with the perturbed SGD method). In our future work, we plan to go further in exploring this relationship.

### v2

@@ 4. Error of the randomized Euler scheme
In this section, we introduce a randomized Euler scheme for the strong approximation of solutions of (1) and investigate its $p$-th moments, the algorithm’s informational cost, and error upper bounds.
If the explicit formula of $a$ is known, the standard Euler-Maruyama scheme can be leveraged to approximate the value of $X(T)$. The scheme is defined as
@@
Its $L^{p}(\Omega)$-error, for $p\geq 2,$ is proportional to $\mathcal{O}(n^{-1/2})$; see Theorem 10.2.2 in [kloeden]. By the standard Euler-Maruyama algorithm, we mean the evaluation of $X_{n}^{E}(T)$ via the standard Euler-Maruyama scheme. In this paper, we define the complexity of an algorithm in terms of the information-based complexity framework; see [ibcbook]. By the informational cost of $X_{n}^{E}(T)$, we mean the total number of scalar evaluations of $a,b$ and $W$. Hence, the informational cost of the standard Euler-Maruyama algorithm is $\Theta(n).$
In the next part, we introduce a randomized Euler algorithm that can only access the information about $H$ in case $a$ is unknown. Let $n,M\in\mathbb{N},h=T/n$ and $t_{k}=kh$ for $k=0,1.\ldots,n.$ randomized Euler scheme is defined as
$\begin{cases}X_{n,M}^{RE}(0)=\eta,\\ X_{n,M}^{RE}(t_{k+1})=X_{n,M}^{RE}(t_{k})+\frac{h}{M}\sum\limits_{j=1}^{M}H(\xi_{j}^{k+1},X_{n,M}^{RE}(t_{k}))+b(X_{n,M}^{RE}(t_{k}))\Delta W_{k},\\ \quad k=0,1,\ldots,n-1\end{cases}$ (7)
where $(\xi_{j}^{k})_{j,k}$ is $\mu$-distributed i.i.d sequence and $\sigma\Big(\bigcup\limits_{k\geq 0,j\geq 1}\sigma(\xi_{j}^{k})\Big)\perp\!\!\!\perp\sigma\Big(\bigcup\limits_{t\geq 0}\mathcal{F}_{t}^{W}\Big)$ which means the randomization is independent of driving Wiener process. Consequently,
$\sigma\Big(X_{n,M}^{RE}(t_{k})\Big){\perp\!\!\!\perp}\sigma\Big(\xi_{1}^{k+1},\ldots,\xi_{M}^{k+1}\Big)$ for all $k=1,\ldots,n$; see the first part of the proof of Lemma 3. The continuous version of the randomized Euler scheme is defined as
$\tilde{X}_{n,M}^{RE}(t)=\eta+\int\limits_{0}^{t}\sum\limits_{k=0}^{n-1}\frac{1}{M}\sum\limits_{j=1}^{M}H(\xi_{j}^{k+1},\tilde{X}_{n,M}^{RE}(t_{k}))\mathds{1}_{(t_{k},t_{k+1}]}(s)\,{\mathrm{d}}s+\int\limits_{0}^{t}\sum\limits_{k=0}^{n-1}b(\tilde{X}_{n,M}^{RE}(t_{k}))\mathds{1}_{(t_{k},t_{k+1}]}(s)\,{\mathrm{d}}W(s)$
for $t\in[0,T].$ The continuous version of standard Euler scheme $\tilde{X}_{n}^{E}$ is defined analogously. From mathematical induction, one obtains that
@@
for all $n$ and $M.$ Furthermore, from fact 2, the randomized Euler scheme takes finite values. Based on this auxiliary observation, the following Lemma can be formulated.
@@
There exists constant $C\in[0,+\infty)$ depending only on $\eta,T$ and parameters of the class $\mathcal{F}(L,K,p),$ such that for all $(H,b)\in\mathcal{F}(L,K,P),M,n\in\mathbb{N}$ the following condition holds
@@
By the randomized Euler algorithm, we mean the evaluation of $X_{n,M}^{RE}(T)$ via the randomized Euler scheme. Similarly, we define the complexity of the algorithm in terms of the information-based complexity framework. The algorithm evaluates the value of $W$ at $n$ points, $b(\cdot)$ at $n$ points and $H(\cdot,\cdot)$ at $nM$ points. Thus, $\operatorname{cost}(X_{n,M}^{RE})=2n+nM=\Theta(nM)$. One may notice that the informational cost differs from the informational cost of the standard (randomized) Euler scheme.
@@
There exists constant $C\in[0,+\infty)$ depending only on $\eta,T$ and parameters of the class $\mathcal{F}(L,K,p),$ such that for all $(H,b)\in\mathcal{F}(L,K,p)$ and $n,M\in\mathbb{N}$ the following inequality holds
@@
In this proof, we leverage a continuous versions of Euler schemes. First, note that
@@
$\begin{split}A_{1,n}^{M}(t)&\leq\|L\|_{1}^{p}\sum\limits_{k=0}^{n-1}\int\limits_{0}^{t}\mathbb{E}\Big\|\tilde{X}_{n,M}^{RE}(t_{k})-\tilde{X}_{n}^{E}(t_{k})\Big\|^{p}\mathds{1}_{(t_{k},t_{k+1}]}(s)\,{\mathrm{d}}s\\ &\leq\|L\|_{1}^{p}\int\limits_{0}^{t}\sup\limits_{0\leq u\leq s}\mathbb{E}\Big\|\tilde{X}_{n,M}^{RE}(u)-\tilde{X}_{n}^{E}(u)\Big\|^{p}\,{\mathrm{d}}s\end{split}$
@@
[…]
@@
Since function function $0,T]\ni t\mapsto\sup\limits_{0\leq u\leq t}\mathbb{E}\|\tilde{X}_{n,M}^{RE}(u)-\tilde{X}_{n}^{E}(u)\|^{p}$ is Borel-measurable (as a non-decreasing function) and bounded (by fact 2), applying the Gronwall’s lemma yields
@@
From Minkowski inequality, Theorem 10.2.2 in [kloeden] and lemma (2), one obtains that
@@ 5. Numerical experiments
This section compares the obtained theoretical results with the outputs of performed simulations. In subsection 5.1, we focus on validating the main results from theorem (1). We estimate $L^{2}(\Omega)$ error of the randomized Euler algorithm and compare it with its informational cost. In subsection 5.2, we inspect randomized Euler algorithm trajectories in the optimization problem of finding local minima. As a byproduct, we show that the method can be used to search for the local minima. We decided to search for the local minima of functions of two variables since such functions can be plotted conveniently. Next, in subsection 5.3, we provide a detailed comparison with other optimization algorithms (optimizers), including Nesterov Accelerated Gradient (NAG), Adaptive Gradient (AdaGrad) and Adaptive Moment Estimation (ADAM). Finally, in subsection 5.4, we provide implementation details in CUDA C. We perform a Monte Carlo simulation with independent samples from the Euler algorithm to estimate the $L^{2}(\Omega)$-error. Hence, to speed up the calculations we leverage CUDA C and parallelize the generation of random samples by running the randomized Euler algorithm on Nvidia Titan V (VG100) GPU under CUDA 10.2. The remaining part of the computations is performed on a single CPU core of Intel(R) Xeon(R) CPU E5-2650 v4 @ 2.20GHz. The implementation utilizes $100$ CUDA blocks and $32$ threads per block.
@@ 5.1. Upper error bound vs informational cost
$b(x_{1},x_{2})=\begin{pmatrix}0.16x_{1}&0.24x_{2}\\ 0.24x_{1}&0.32x_{2}\\ \end{pmatrix}$.
@@
The figure 1 shows results obtained via numerical experiments for $\varepsilon_{6400}(X^{RE}_{n,n}(1))$ and various values of $n.$ Note that, according to the theoretical findings from before, the informational cost of the algorithm is of order $\operatorname{cost}(X^{RE}_{n,n}(1))=\Theta(n^{2})$, and its $L^{2}(\Omega)$-error is of order $\|X(1)-X^{RE}_{n,n}(1)\|_{2}=\mathcal{O}(n^{-1/2})$. Hence, $\operatorname{cost}(X^{RE}_{n,n}(1))=\mathcal{O}(\|X(1)-X_{n,n}^{RE}(1)\|_{2}^{-4})$. Similarly, in numerical experiments, we obtain a strong linear correlation between estimated error and cost. It may suggest that in this case, the theoretical findings on the relation between error and cost might be improved and asymptotically equal in terms of $\Theta$ rather than $\mathcal{O}$. Nonetheless, this cannot be addressed without any theoretical findings on error lower bounds which we leave as an open question for future research. Finally, it is worth noting that the slope coefficient is equal to 3.82-3.82 which is also seemingly close to 4.-4.
@@ 5.2. Optimization with randomized Euler algorithm
5.3. Comparison with other optimization methods
In this subsection, we consider $5$ different optimization algorithms/methods, also known as optimizers. In particular, SGD and PSGD optimizers are implemented in line with the Euler scheme definition from section 4, where PSGD relies on $b(x_{1},x_{2})=0.3I_{2}$. From the previous subsection, we know that the number of samples can make a real impact on the cost function formula, and henceforth, the optimization method’s trajectory. To have reliable comparison results i.e., undistrupted by the impact of the number of samples, we only consider the following deterministic cost functions to be optimized:
  1. (i)
    paraboloid function
    $f(x_{1},x_{2})=x_{1}^{2}+x_{2}^{2},$
  2. (ii)
    Himmelblau’s function
    $f(x_{1},x_{2})=(x_{1}^{2}+x_{2}-11)^{2}+(x_{1}+x_{2}^{2}-7)^{2},$
  3. (iii)
    sigmoid-activation loss function
    $f(x_{1},x_{2})=[0.5-(1+\exp(-(x_{1}+x_{2})))^{-1}]^{2}.$
Figure 3. Trajectories of various optimizers searching local minima.
For instance, in figures 3-5, we can find paths of various optimizers for given optimization problems where all the optimizers start at $4$ common but randomly selected initial points. Next, we examine optimization errors, which are measured in terms of distance from current point to its closest local minimum at every single iteration. To have a better insight into paths presented in the previous figures, in figures 6-8, we can find a comparison of errors for one of the previously selected initial points. Next, in numerical experiments, we took the following quantitative approach. For each optimization problem, we generated $1000$ random initial values where the distance from each value to its closest local minimum varied between $1.5$ and $3$. Then mean errors at the log-scale were computed for each optimizer respectively, which can be found in figures 9-11. Finally, based on such values, we also computed the quantile spread of log-scale errors for the optimization algorithms that can be found in figures 12-14. From the obtained results, we see that the PSGD algorithm works well. Nevertheless, due to its random corrections (in order to avoid plateaus), it is not fully able to find local minima, and henceforth it gets stuck at a certain error level, which can be easily beaten by adaptive methods like ADAM or NAG. In the end, we suspect that a better-designed diffusion function $b$ may lead to finer results, which we leave as an open question.
Figure 4. Trajectories of various optimizers searching local minima.
Figure 5. Trajectories of various optimizers searching local minima.
Figure 6. Errors of various optimizers starting at the selected point.
Figure 7. Errors of various optimizers starting at the selected point.
Figure 8. Errors of various optimizers starting at the selected point.
Figure 9. Mean errors of various optimizers starting at randomly selected points.
Figure 10. Mean errors of various optimizers starting at randomly selected points.
Figure 11. Mean errors of various optimizers starting at randomly selected points.
Figure 12. Quantile spread for errors of various optimizers starting at randomly selected points.
Figure 13. Quantile spread for errors of various optimizers starting at randomly selected points.
Figure 14. Quantile spread for errors of various optimizers starting at randomly selected points.
@@ 5.4. Implementation details
In this subsection, we provide implementation details in CUDA C.
1template <typename type>
2__device__ type func_H(int d, type t, type* x);
3template <typename type>
4__device__ type func_b(int d, int k, type t, type* x);
5
6template <typename type>
7void euler_single_step(type t, type* X_prev, int X_dim, type h, type* ksi, int ksi_num, type* dW, int dW_dim, type* X) {
8 // (1)
9 for (int d = 0; d < X_dim; d++){
10 // (2)
11 for (int j = 0; j < ksi_num; j++) {
12 X[d] += func_H<type>(d, ksi[j], X_prev) * (h/ksi_num);
13 }
14 // (3)
15 for (int k = 0; k < dW_dim; k++) {
16 X[d] += func_b<type>(d, k + 1, t, X_prev) * dW[k];
17 }
18 }
19}
Listing 1: Numerical scheme implementation
@@ 16 X[d] += func_b<type>(d, k + 1, t, X_prev) * dW[k];
    Traverse through system of $d$ equations.
@@
    Add drift component.
@@
    Add diffusion component.
1#define FP float
2
3extern struct curandState_t;
4void* malloc(size_t size);
5void* memset(void* dest, int ch, size_t count);
6void* memcpy(void * destination, const void * source, size_t num);
7void free(void *ptr);
8__device__ int LCM(int x, int y);
9__device__ float curand_normal(curandState_t* state);
10__device__ double sqrt(double x);
11__device__ float curand_uniform(curandState_t* state);
12template <typename type>
13__device__ type drift_inverse_cdf(type x);
14
15__device__ void sample_from_coupled_X(int ML, int Ml, int nL, int nl, int dW_dim, FP* x0, FP T, FP* XL, FP* Xl, int X_dim, curandState_t* state) {
@@ 35 FP t, dW;
52 ksi[j] = drift_inverse_cdf<FP>((FP) curand_uniform(state)); // (7)
@@ 6. Conclusions
We have investigated the error of the randomized Euler scheme in the case when the drift of underlying SDE has a special integral form. As a result, we proved a theorem that describes the error upper bound of the algorithm in terms of the algorithm parameters i.e., grid density and the number of summands in the approximation of the drift. It turns out that various error upper bounds can be found depending on the argument dependence of drift’s integrand. Moreover, we suspect that the informational cost of the algorithm could further be reduced with the usage of the Multilevel Monte Carlo algorithm regarding the approximation of the drift. Nevertheless, we leave this here as an open question for future research. In the remaining part of the paper, we conducted numerical experiments using GPU architecture. The numerical experiments show that the estimated error conforms with the obtained upper bounds. Hence, we show that the randomized Euler scheme is a suitable method for approximating trajectories of SDEs (1). By product, in numerical experiments, we show that the algorithm can be used as an optimization method (due to its connection with the perturbed SGD method). Nevertheless, there are no strict theoretical results on finding the actual local minima. In our future work, we especially plan to go further in exploring the relationship between SDEs and gradient descent algorithms.